OpenStudy (jhulian):

got stuck... 2sin(squared) theta =1

OpenStudy (kawii2004):

okay

OpenStudy (jhulian):

help me solve

OpenStudy (kawii2004):

umm wait

OpenStudy (anonymous):

Do you mean 2sin^2theta=1

OpenStudy (jhulian):

yes

OpenStudy (anonymous):

\[2 \sin ^2\theta=1,or~1-\cos 2\theta=1\] \[\cos 2\theta=0=\cos (2n+1)\frac{ \pi }{ 2 }\] \[2 \theta=\left( 2n+1 \right)\frac{ \pi }{ 2 },\theta=\left( 2n+1 \right)\frac{ \pi }{ 4 }\]

OpenStudy (anonymous):

here n is an integer.

OpenStudy (jhulian):

thanks

OpenStudy (anonymous):

2sin²(θ) = 1 sin²(θ) = 1/2 sin(θ) = ± 1/√2 sin(θ) = ± (√2)/2 First case: sin(θ) = + (√2)/2 θ = π/4 θ = π - (π/4) = 3π/4 Second case: sin(θ) = - (√2)/2 θ = 2π - (π/4) = 7π/4 θ = π + (π/4) = 5π/4 → Solution = { π/4 ; 3π/4 ; 5π/4 ; 7π/4 }

OpenStudy (jhulian):

Thanks princess!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

You are quite welcome @Jhulian if you no longer have any questions about this subject please do close the question. Thank you very much.(: